7. Tempo Electronics, Inc., has an inventory of 5,000 unique electronic chip originally purchased at $2.50 each; their market value is now $ 5 each. The production department has proposed to use these by putting each one together with $6 worth of labor and other materials to produce a wristwatch that would be sold for $10. Should that proposal be implemented? Explain from the viewpoints of economic profit and opportunity costs.
1
Expert's answer
2020-11-05T09:23:25-0500

1.Let's calculate the profit in the first case:

"5\times5000-2.50\times5000=25000-12500=12500"

2.Let's calculate the profit in the second case:

"10\times5000-6\times5000-2.5\times5000=7500"

opportunity costs: 30 000(5000*6)

economic profit: 7500

Economic profit is less than profit in the first case, so it is not worth going along the second production path, or then it is necessary to increase the sales price of the finished product